The weights (in grams) of the contents of several small bottles are 4, 3, 5, 4, 5 and

3. Find mean, Standard deviation and the coefficient of variation



1
Expert's answer
2021-08-09T16:54:19-0400

Mean: "\mu=\frac{4+3+5+4+5+3}{6}=\frac{24}{6}=4."

Variance: "\sigma^2=\frac{(4-4)^2+(3-4)^2+(5-4)^2+(4-4)^2+(5-4)^2+(3-4)^2}{6}=\frac{4}{6}\approx0.6667."

Standard deviation: "\sigma=\sqrt\frac{2}{3}\approx0.8165."

Coefficient of variation: "CV=\frac{\sigma}{\mu}=\frac{0.8165}{4}=0.2041=20.41\%."